OpenStudy (anonymous):

1. Complete the table below to solve the equation 2.5x – 10.5 = 64(0.5x). x f(x) = 2.5x – 10.5 g(x) = 64(0.5x) 2 3 4 5 6

OpenStudy (anonymous):

@Callisto @mathslover @uri @Preetha

OpenStudy (anonymous):

in need of help,

OpenStudy (anonymous):

@Koikkara @Warriorz13

OpenStudy (anonymous):

i thought this was suppose to be a helping sight.

OpenStudy (warriorz13):

-__-

OpenStudy (anonymous):

?

OpenStudy (anonymous):

please help.

OpenStudy (anonymous):

i have more questions if you want me to post them.

OpenStudy (anonymous):

PLEASE HELP (:

OpenStudy (anonymous):

@Warriorz13 @Koikkara

OpenStudy (koikkara):

Hmm... I have no idea what they want from you, but I'm just going to solve that equation: 2.5x – 10.5 = 64(0.5x) 2.5x - 10.5 = 32x 2.5x - 32x = 10.5 -29.5x = 10.5 x = -10.5/29.5 Does this help ?

OpenStudy (anonymous):

yes thank you.

OpenStudy (koikkara):

@alyssa_fields50 Thank U !! Keep In Touch, with Open study !!

OpenStudy (anonymous):

i could use alittle more help if you wouldnt mind @Koikkara

OpenStudy (koikkara):

sure, i will try...!

OpenStudy (anonymous):

A wooded area in a state park has a mixture of different types of trees. There are 800 pine trees and 50 oak trees. The number of pine trees is decreasing at a rate of 5% per year. The number of oak trees is increasing at a rate of 15% per year. If these trends continue: Write two functions to model this situation and graph those two functions on the same coordinate grid. During what year in the future will the park have approximately the same number of pine and oak trees? How many of each type of tree will there be at that time?

OpenStudy (anonymous):

are you still working on this ?

OpenStudy (anonymous):

@Koikkara

OpenStudy (koikkara):

lol... i was out in bio.... just a min, let me have a look !

OpenStudy (anonymous):

thank youu.

OpenStudy (anonymous):

have you figured anything out.

OpenStudy (koikkara):

http://openstudy.com/study#/updates/5258df2fe4b0ede9e4485087 @alyssa_fields50 better look at it ! because i don't know how to explain ! Good day to you !